The Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 10,000 $ 15,000 $ 25,000
Estimated variable manufacturing overhead per machine-hour $ 1.40 $ 2.20
Job P Job Q
Direct materials $ 13,000 $ 8,000
Direct labor cost $ 21,000 $ 7,500
Actual machine-hours used:
Molding 1,700 800
Fabrication 600 900
Total 2,300 1,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)

12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.)

13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations.)

15. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1
Predetermined overhead rates:
Molding Department 5.40 =1.4+(10000/2500)
Fabrication Department 12.20 =2.2+(15000/1500)
Job P Job Q
Direct materials 13000 8000
Direct labor cost 21000 7500
Manufacturing overhead applied 16500 15300
Total manufacturing cost 50500 30800
11
Manufacturing overhead applied:
Job P 7320 =600*12.20
Job Q 10980 =900*12.20
12
Total manufacturing cost 50500
Divide by units 20
Unit product cost 2525
13
Total manufacturing cost 30800
Divide by units 30
Unit product cost 1027
14
Job P Job Q
Total manufacturing cost 50500 30800
Add: Markup @ 80% 40400 24640
Total price for the job 90900 55440
Divide by units 20 30
Selling price per unit 4545 1848
15
Cost of goods sold for March 81300 =50500+30800